### v1

# OminiControl: Minimal and Universal Control for Diffusion Transformer

Abstract
In this paper, we introduce OminiControl, a highly versatile and parameter-efficient framework that integrates image conditions into pre-trained Diffusion Transformer (DiT) models. At its core, OminiControl leverages a parameter reuse mechanism, enabling the DiT to encode image conditions using itself as a powerful backbone and process them with its flexible multi-modal attention processors. Unlike existing methods, which rely heavily on additional encoder modules with complex architectures, OminiControl (1) effectively and efficiently incorporates injected image conditions with only 0.10.1~{}0.10.1additional parameters, and (2) addresses a wide range of image conditioning tasks in a unified manner, including subject-driven generation and spatially-aligned conditions such as edges, depth, and more. Remarkably, these capabilities are achieved by training on images generated by the DiT itself, which is particularly beneficial for subject-driven generation. Extensive evaluations demonstrate that OminiControl outperforms existing UNet-based and DiT-adapted models in both subject-driven and spatially-aligned conditional generation. Additionally, we release our training dataset, Subjects200K, a diverse collection of over 200,000 identity-consistent images, along with an efficient data synthesis pipeline to advance research in subject-consistent generation.  [^1] Code and dataset are available at https://github.com/Yuanshi9815/OminiControl
Figure 1: Results of our OminiControl on both subject-driven generation (top) and spatially-aligned tasks (bottom). The small images in red boxes show the input conditions.
1 Introduction
Diffusion models[9, 28, 25] have revolutionized the field of visual generation, demonstrating remarkable capabilities that significantly outperform traditional approaches like Generative Adversarial Networks (GANs)[6] in terms of image quality and diversity. While these models excel at generating highly realistic imagery, a critical challenge persists: enabling precise and flexible control over the generation process to accommodate diverse and complex user requirements.
Text-based conditioning has been a cornerstone in advancing controllable generation[28, 25, 13, 2, 23, 35], offering an intuitive interface for users to specify their desired outputs. However, text prompts alone often fail to convey precise spatial details and structural attributes that users wish to control. Consequently, recent research has explored complementary conditioning modalities for guiding diffusion models, with image-based control emerging as a particularly effective approach[41, 39, 22, 43, 15]. This multi-modal conditioning strategy enables more detailed and accurate control over the generation process, addressing the limitations inherent in purely text-based interfaces.
Current image conditioning methods can be broadly categorized into spatially aligned and non-spatially aligned approaches. Spatially aligned tasks such as sketch-to-image and inpainting require direct correspondence between conditioning and output images, typically achieved through methods like ControlNet[41] that inject conditioning features in a spatially-preserving manner. In contrast, non-spatially aligned applications including subject driven generation and style transfer, as demonstrated by IP-Adapter[39], often employ pre-trained encoders like CLIP[27] to extract global features for integration via cross-attention mechanisms.
Despite the effectiveness of existing image-conditioned approaches, they present several limitations that hinder their efficiency and flexibility [41, 39, 22]. First, most existing methods are designed specifically for UNet-based architectures [29, 41, 22, 39, 31, 42, 43, 40, 24, 16], as seen in Stable Diffusion models [28, 25]. While these approaches work well with UNet’s encoder-decoder structure, they may not translate effectively to the more advanced Diffusion Transformer (DiT) models[23] that have demonstrated superior image generation quality[13, 2]. Additionally, current approaches typically specialize in either spatially aligned [22, 43, 41] or non-spatially aligned tasks [39, 42, 17, 15, 12], lacking a unified architecture to handle both control types effectively. This specialization often requires practitioners to employ different methods for different control scenarios, increasing system complexity and implementation overhead. Furthermore, these methods rely heavily on additional network structures [39, 42, 22, 43, 41, 17], which introduce substantial parameter overhead.
To address these limitations, we propose a parameter-efficient approach for incorporating image-based control into DiT architectures[23]. Our method reuse the model’s existing VAE encoder[28] to process conditioning images. Following the same token processing pipeline as noisy image tokens, we augment the encoded features with learnable position embeddings[34], and seamlessly integrate them alongside latent noise in the denoising network. This design enables direct multi-modal attention interactions[23, 30] between condition and generation tokens throughout the DiT’s transformer blocks, facilitating efficient information exchange and control signal propagation.
Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
We implemented our method on the high-performing DiT-structured diffusion model, FLUX.1-dev[13], a large-scale model containing 12 billion parameters. Extensive experiments on edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation indicate that our DiT-based approach yields better results compared to both UNet-based implementations[41, 39, 7] and their community adaptations on the FLUX.1 model [14, 37].
For identity-preserving generation, we developed a novel data synthesis pipeline that generates high-quality, identity-consistent image pairs. Using this pipeline, we created a comprehensive dataset comprising over 200,000 diverse images. To facilitate future research in this direction, we will release both our dataset and the complete pipeline implementation as open-source resources [^2] More details are available in the supplementary material..
In summary, we highlight our contributions as follows:
  1. 1.
    We present a parameter-efficient method for enabling image-conditioned control in Diffusion Transformer (DiT) models, achieving both spatially aligned and non-spatially aligned control within a unified framework.
  2. 2.
    We demonstrate the effectiveness of our approach through extensive experiments across diverse control tasks, including edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation, consistently outperforming existing methods on both UNet implementations and their DiT adaptations.
  3. 3.
    We develop and release Subjects200K, a high-quality dataset of over 200,000 subject-consistent images, along with an efficient data synthesis pipeline, providing valuable resources to the research community for further exploration of subject-consistent generation tasks.
2 Related works
2.1 Diffusion-based models
Diffusion-based methods have emerged as a powerful framework for image generation[9, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1]. Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28]. To further enhance generative capabilities, large-scale transformer architectures have been integrated into these frameworks, leading to advanced models like DiT[23, 2, 13, 3]. Building on these architectural innovations, FLUX[13] incorporates transformer-based design with flow matching objectives [18], achieving state-of-the-art generation performance.
2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [41], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition information to the denoising network’s hidden states, inherently limiting their effectiveness for non-spatial tasks like subject-driven generation. IP-Adapter [39] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [42] further enhances identity preservation in image-conditioned tasks. Despite these advances [15, 19, 5], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
3 Methods
3.1 Preliminary
The Diffusion Transformer (DiT) model [23], employed in architectures like FLUX.1 [13], Stable Diffusion 3 [28], and PixArt [2], uses a denoising network of transformer blocks to refine noisy image tokens iteratively.
Each transformer block processes two types of tokens: noisy image tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). These tokens are embedded into hidden states $X$ and $C_{\text{T}}$, which maintain consistent shapes throughout the transformer blocks.
In each DiT block, after normalizing $X$ and $C_{\text{T}}$, they are processed by the core MM-Attention module [30], which employs Rotary Position Embedding (RoPE) [34] to incorporate positional dependencies across tokens. For a token at position $(i,j)$ in the 2D grid, RoPE applies rotation matrices to the query and key projections:
$Q_{X}(i,j)$ $=W_{Q}\left(X_{i,j}\cdot R(i,j)\right),$ (1)
$K_{X}(i,j)$ $=W_{K}\left(X_{i,j}\cdot R(i,j)\right),$ (2)
where $R(i,j)$ is the rotation matrix at position $(i,j)$. Similarly, the text condition tokens $C_{\text{T}}$ have their query and key projections defined in the same way, with all text token positions set to $(0,0)$ in FLUX.1.
After applying RoPE, the queries, keys, and values from both token types are concatenated to form unified matrices $Q_{Z}$, $K_{Z}$, and $V_{Z}$, representing the combined token set $Z=[X;C_{\text{T}}]$. The MM-Attention operation is then computed as:
$\text{MMAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}\right)V_{Z},$ (3)
enabling interactions between image and text condition tokens through the attention mechanism.
Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
3.2 Image condition integration
Our approach first encodes the condition image through the model’s VAE, projecting it into the same latent space as the noisy image tokens to form $\mathbf{C}_{\text{I}}\in\mathbb{R}^{N\times d}$.
Previous methods like ControlNet [41] and T2I-Adapter [22] incorporate the condition image by spatially aligning and adding its hidden states directly to those of the noisy image tokens:
$H_{X}=H_{X}+H_{C_{\text{I}}},$ (4)
where $H_{X}$ represents the combined hidden states for further processing, with $H_{C_{\text{I}}}$ being the hidden states from the condition image. While this approach proves effective for spatially aligned tasks, it faces two key limitations: (1) it lacks flexibility when handling non-aligned scenarios, and (2) even in spatially aligned cases, the direct addition of hidden states constrains token interactions, potentially limiting the model’s performance.
(a) Training losses for different image condition integration methods.
(b) Training loss for shared vs. shifting position.
Figure 4: Training loss comparisons.
In contrast, to enable non-aligned control tasks and provide greater conditioning flexibility, our method processes condition image tokens uniformly with text and noisy image tokens, integrating them into a unified sequence:
$Z=[X;C_{\text{T}};C_{\text{I}}],$ (5)
where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$. This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2). rect addition approach.
Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$. Strong diagonal patterns indicate effective spatial alignment. (b) Subject-driven generation task, with input condition and output image. Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
Comparative analysis shows that our approach achieves better results than the direct additive method, as demonstrated by the attention visualization in Figure 3. Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the di The effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
3.3 Adaptive position embedding
The integration of condition images requires careful consideration of positional information to ensure effective interaction between condition and target images. Previous methods often assume rigid spatial alignment between condition and output images, limiting their applicability to non-aligned tasks. Moreover, the relative positioning between condition and target tokens can significantly impact the model’s learning efficiency and generalization ability. Leveraging our unified sequence design where condition tokens are concatenated with other tokens, we have the flexibility to explore different positional encoding strategies.
In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information. For a 512$\times$512 input image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token receives a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$. This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the original image. However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)). Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
3.4 Condition strength factor
The unified attention mechanism we adopt not only enables flexible token interaction but also allows us to precisely control the influence of condition images. Specifically, we designed a method that allows for manual adjustment of the condition image’s effect during inference. For a given strength factor $\gamma$, setting $\gamma=0$ removes the condition image’s influence, resulting in an output based purely on the original input. At $\gamma=1$, the output fully reflects the condition image, and as $\gamma$ increases beyond 1, the condition’s effect becomes even more pronounced.
To achieve this controllability, we introduce a bias term into the original MM-Attention operation. Specifically, we modify Equation 3 to:
$\text{BiasedAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}+\text{bias}(\gamma)\right)V_{Z},$ (6)
[…]
$\text{bias}(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (7)
This design ensures that the strength factor $\gamma$ only affects the attention weights between noisy image tokens and condition image tokens, while maintaining the original attention patterns for text tokens and within-modality interactions.
3.5 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, vehicles, and animals, totaling over 200,000 images. This dataset, along with the generation pipeline, will be publicly released.
Training models for subject-consistent generation typically requires paired images that maintain identity consistency while exhibiting variations in pose, lighting, and other attributes. Previous methods like IP-Adapter [39] use identical images for conditioning and target pairs, which proves effective for their approaches. However, in our framework, this setup leads to overfitting, causing the model to generate outputs nearly identical to the inputs.
To overcome these limitations, we developed a dataset featuring images that preserve subject identity while incorporating natural variations. While existing datasets [31, 12, 15, 17] address similar needs, they often face constraints in either quality or scale. We therefore propose a novel synthesis pipeline leveraging FLUX’s inherent capability to generate pairs of visually related images from carefully crafted prompts.
Our pipeline utilizes ChatGPT-4o to generate over 20,000 diverse image descriptions, which guide FLUX in producing more than 200,000 images (Figure 6). The generated images undergo quality assessment using ChatGPT-4o’s vision capabilities, ensuring both consistency and diversity in the final dataset. The curated dataset and complete generation pipeline are publicly available [^3] More details are provided in the supplementary material. Dataset and code are available at https://github.com/Yuanshi9815/Subjects200K.
4 Experiment
4.1 Setup
Tasks and base model. We evaluate our method on two categories of conditional generation tasks: spatially aligned tasks (including Canny-to-image, depth-to-image, masked-based inpainting, and colorization) and subject-driven generation. We build our method upon FLUX.1 [13], a latent rectified flow transformer model for image generation. By default, we use FLUX.1-dev to generate images for spatially aligned tasks. In subject-driven generation tasks, we switch to FLUX.1-schnell as we observed it tend to produce better visual quality.
Condition Model Method Controllability General Quality Text Consistency
[…]
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.75
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.74
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.53
Ours 0.38 20.63 0.40 0.61 75.91 0.76
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.726
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.72
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.54
Ours 903 27.26 0.39 0.55 75.06 0.728
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.78
Ours 132 11.49 0.87 0.39 67.63 0.87
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.85
Ours 24 10.23 0.73 0.43 70.74 0.90
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.84
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.80
Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks. We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score). For F1-Score, higher is better; for MSE, lower is better. Best results are shown in bold.
Implement details. Our method utilizes LoRA [4]for fine-tuning the base model with a default rank of 4. To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 when processing non-condition tokens by default.
Training. Our model is trained with a batch size of 1 and gradient accumulation over 8 steps (effective batch size of 8). We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01. The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each). For spatially aligned tasks, models are trained for 50,000 iterations, while subject-driven generation models are trained for 15,000 iterations.
Baselines. For spatially-aligned tasks, we compare our method with both the original ControlNet [41] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [14], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [39], evaluating its implementations FLUX.1 [37].
Evaluation metrics. We evaluate our model on both spatially aligned tasks and subject-driven generation. For spatially aligned tasks, we assess two aspects: generation quality and controllability. Generation quality is measured using FID [8], SSIM, MAN-IQA [38], and MUSIQ [10] for visual fidelity, along with CLIP Score [27] for semantic consistency. For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.). For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation. Detailed evaluation methodology is presented in Appendix B.1.
Evaluation protocol. We conducted evaluations on two datasets. For spatially-aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
4.2 Main result
Figure 7: Qualitative results comparing different methods. Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization. Right: Subject-driven generation with beverage can, shoes and robot toy. Our method demonstrates superior controllability and visual quality across all tasks.
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on Canny-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [38], and MUSIQ [10] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [8] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines. Our method demonstrates superior performance, particularly in identity preservation and modification accuracy. Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%. The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing. These quantitative results are further corroborated by user studies presented in Appendix B.1.
Comparative paramter efficiency. As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods. For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%). Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
Methods Base model Parameters Ratio
ControlNet SD1.5 / 860M 361M $\sim$42%
T2I-Adapter 77M $\sim$9.0%
IP-Adapter 449M $\sim$52.2%
ControlNet FLUX.1 / 12B 3.3B $\sim$27.5%
IP-Adapter 918M $\sim$7.6%
Ours FLUX.1 / 12B 14.5M / 48.7M w/ Encoder $\sim$0.1/ $\sim$0.4w/ Encoder
Table 2: Additional parameters introduced by different image conditioning methods. For IP-Adapter, the parameter count includes the CLIP Image encoder. For our method, we also report results when using the original VAE encoder from FLUX.1.
4.3 Empirical studies
Effect of training data.
Figure 9: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
For subject-driven generation, our model takes a reference image of a subject (e.g., a plush toy or an object) and a text description as input, aiming to generate novel images of the same subject following the text guidance while preserving its key characteristics.
To validate the effectiveness of our Subjects200K dataset described in Section 3.5, we compare two training strategies for this task. The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images. The second approach utilizes our Subjects200K dataset. As shown in Figure 9, the model trained with data augmentation only learns to replicate the input conditions with minimal changes. In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose. Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction. In contrast, our Subjects200K-trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like canny-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
Figure 10: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
Impact of LoRA rank. We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task. As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects: image quality (measured by FID and SSIM), condition control capability (measured by F1 Score), while maintaining competitive text-image consistency (measured by CLIP-Score). Notably, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment where it achieves the highest CLIP-Score of 0.765, showing the efficiency of our approach even with limited parameters.
Study Setting FID $\downarrow$ SSIM $\uparrow$ F1 Score $\uparrow$ CLIP Score $\uparrow$
LoRA Rank 1 21.09 0.412 0.385 0.765
2 21.28 0.411 0.377 0.751
4 20.63 0.407 0.380 0.761
8 21.40 0.404 0.3881 0.761
16 19.71 0.425 0.407 0.764
Condition Blocks Early 25.66 0.369 0.23 0.72
Full 20.63 0.407 0.38 0.76
Table 3: Ablation studies on (1) LoRA rank for the Canny-to-image task and (2) condition signal integration approaches. Results show that LoRA rank of 16 and full-depth integration achieve the best performance. Rows with blue background indicate our default settings (LoRA rank=4, Full condition integration). Best results are in bold.
Conditioning depth. FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens. As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability over the generation process. This suggests that allowing the condition signals to influence the entire transformer stack is crucial for achieving the desired levels of control over the output. Notably, this finding indicates that the preview approaches[14, 37, 39, 41, 22] of inserting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
5 Conclusion
OmniControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OmniControl’s scalability and effectiveness in diffusion models.
Supplementary Material
Appendix A Details of Subjects200K datasets
We present a comprehensive synthetic dataset constructed to address the limitations in scale and image quality found in previous datasets [31, 12, 15, 17]. Our approach leverages FLUX.1-dev [13] to generate high-quality, consistent images of the same subject under various conditions.
Subjects200K dataset currently consists of two splits, both generated using similar pipelines. Split-1 contains paired images of objects in different scenes, while Split-2 pairs each object’s scene image with its corresponding studio photograph. Due to their methodological similarities, we primarily focus on describing the synthesis process and details of Split-2, although both splits are publicly available. Our complete Subjects200K dataset can be fully accessed via this link.
A.1 Generation pipeline
Our dataset generation process consists of three main stages: description generation, image synthesis, and quality assessment.
Description Generation We employed ChatGPT-4o to create a hierarchical structure of descriptions: We first generated 42 diverse object categories, including furniture, vehicles, electronics, clothing, and others. For each category, we created multiple object instances, totaling 4,696 unique objects. Each object entry consists of: (1) A brief description, (2) Eight diverse scene descriptions, (3) One studio photo description. Figure S2 shows a representative example of our structured description format.
Image Synthesis We designed a prompt template to leverage FLUX’s capability of generating paired images containing the same subject. Our template synthesizes a comprehensive prompt by combining a brief object description with two distinct scene descriptions, ensuring subject consistency while introducing environmental variations.
The detailed prompt structure is illustrated in Figure S3. For each prompt, we set the image dimensions to 1056×528 pixels and generated five images using different random seeds to ensure diversity in our dataset. During the training process, we first split the paired images horizontally, then performed central cropping to obtain 512×512 pixel image pairs. This padding strategy was implemented to address cases where the generated images were not precisely bisected, preventing potential artifacts from appearing in the wrong half of the split images.
Figure S1: Examples of successful and failed generation results from Subjects200K dataset. Green checks indicate successful cases where subject identity and characteristics are well preserved, while red crosses show failure cases.
Quality assessment We leveraged ChatGPT-4o’s vision capabilities to rigorously evaluate the quality of images generated by FLUX.1-dev. The assessment focused on multiple critical aspects:
  • Image composition: Verifying that each image properly contains two side-by-side views.
  • Subject consistency: Ensuring the subject maintains identity across both views.
  • Image quality: Confirming high resolution and visual fidelity.
To maintain stringent quality standards, each image underwent five independent evaluations by ChatGPT-4o. Only images that passed all five evaluations were included in our training dataset. Figure S1 presents representative examples from our quality-controlled dataset.
"brief_description":
"A finely-crafted wooden seating piece.",
"scene_descriptions": [
"Set on a sandy shore at dusk, it faces the ocean with a gentle breeze rustling nearby palms, bathed in soft, warm twilight.",
"Positioned in a bustling urban cafe, it stands out against exposed brick walls, capturing the midday sun through a wide bay window."
// Additional six scene descriptions omitted
"studio_photo_description":
"In a professional studio against a plain white backdrop, it is captured in three-quarter view under uniform high-key lighting, showcasing the delicate grain and smooth of its finely-crafted surfaces."
Figure S2: An example of our structured description format for dataset generation.
prompt_1 = f"Two side-by-side images of the same object: {brief_description}"
prompt_2 = f"Left: {scene_description1}"
prompt_3 = f"Right: {scene_description2}"
prompt_image = f"{prompt_1}; {prompt_2}; {prompt_3}"
Figure S3: Our prompt template for paired image generation. The template combines a brief object description with two distinct scene descriptions to maintain subject consistency while varying environmental conditions.
A.2 Dataset Statistics
In Split-2, we first generated 42 distinct object categories, from which we created and curated a set of 4,696 detailed object instances. Then we combine these descriptions to generate 211,320 subject-consistent image pairs. Through rigorous quality control using GPT-4o, we selected 111,767 high-quality image pairs for our final dataset. This extensive filtering process ensured the highest standards of image quality and subject consistency, resulting in a collection of 223,534 high-quality training images.
Appendix B Additional experimental results
B.1 Evaluation for subject-driven generation
Framework and criteria. To systematically evaluate subject-driven generation quality, we establish a framework with five criteria assessing both preservation of subject characteristics and accuracy of requested modifications:
  • Identity Preservation: Evaluates preservation of essential identifying features (e.g., logos, brand marks, distinctive patterns)
  • Material Quality: Assesses if material properties and surface characteristics are accurately represented
  • Color Fidelity: Evaluates if colors remain consistent in regions not specified for modification
  • Natural Appearance: Assesses if the generated image appears realistic and coherent
  • Modification Accuracy: Verifies if the changes specified in the text prompt are properly executed
Method Identity Material Color Natural Modification Average
preservation quality fidelity appearance accuracy score
Average over 5 random seeds
IP-Adapter (SD 1.5) 29.4 86.1 45.3 97.9 17.0 55.1
SSR-Encoder 46.0 92.0 54.2 96.3 28.5 63.4
IP-Adapter (FLUX) 11.8 65.8 30.8 98.1 57.7 52.8
Ours 50.6 84.3 55.0 98.5 75.8 72.8
Best score over 5 random seeds
IP-Adapter (SD 1.5) 56.3 98.9 70.1 99.7 37.2 72.5
SSR-Encoder 64.3 99.2 74.4 99.1 53.6 78.1
IP-Adapter (FLUX) 27.5 86.1 53.6 99.9 74.9 68.4
Ours 82.3 98.0 88.4 100.0 90.7 91.9
Table S1: Quantitative evaluation results (in percentage) across different evaluation criteria. Higher values indicate better performance.
Figure S4: User study results comparing different methods across three metrics: identity consistency, text-image alignment, and visual coherence.
User studies. To further validate our approach, we conducted user studies collecting 375 valid responses. Participants evaluated the generated images across three key dimensions: identity consistency, text-image alignment, and visual coherence between subjects and backgrounds. The results shown in Figure S4 corroborate our quantitative findings, with our method achieving superior performance across all evaluation criteria.
B.2 Additional generation results
We showcase more generation results from our method. Figure S5 presents additional results on the DreamBooth dataset, while Figure S6 demonstrates our method’s effectiveness on other subject-driven generation tasks.
Figure S5: More results on Dreambooth dataset.
Figure S6: More results on other subject-driven generation tasks.

### v2

@@
 This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2).
-rect addition approach.
 Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$.
@@
 Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
-Comparative analysis shows that our approach achieves better results than the direct additive method, as demonstrated by the attention visualization in Figure 3.
-Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the di The effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
+The comparative result shows that our approach achieves higher generation quality and better alignment with the conditions compared to the direct adding method, as illustrated in Figure 3.
+Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the direct adding approach.
+Besides, the effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
 3.3 Adaptive position embedding
-The integration of condition images requires careful consideration of positional information to ensure effective interaction between condition and target images.
-Previous methods often assume rigid spatial alignment between condition and output images, limiting their applicability to non-aligned tasks.
-Moreover, the relative positioning between condition and target tokens can significantly impact the model’s learning efficiency and generalization ability.
-Leveraging our unified sequence design where condition tokens are concatenated with other tokens, we have the flexibility to explore different positional encoding strategies.
+Our unified sequence design allows for flexible integration of condition image tokens, but this requires incorporating positional information to ensure effective interaction with noisy image tokens.
+The relative positioning of these tokens is critical, as it directly affects the model’s learning efficiency and generalization capability.
 In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information.
-For a 512$\times$512 input image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token receives a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$.
+For a 512$\times$512 target image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token is assigned a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$.
 This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
-For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the original image.
+For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the noisy image.
 However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)).
@@
 Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks.
-Our method achieves the highest F1-Score of 0.38 on Canny-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14].
+Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14].
 In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [38], and MUSIQ [10] scores.
@@
 Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$.
-Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like canny-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
+Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like depth-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
 Figure 10: Comparison of models trained with different data.
@@
 5 Conclusion
-OmniControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules.
-Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation.
-Results confirm OmniControl’s scalability and effectiveness in diffusion models.
+OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules.
+Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation.
+Results confirm OminiControl’s scalability and effectiveness in diffusion models.
+6 Acknowledgment
+We would like to acknowledge that the computational work involved in this research work is partially supported by NUS IT’s Research Computing group using grant numbers NUSREC-HPC-00001.
 Supplementary Material

### v4

@@
 1 Introduction
-Diffusion models[9, 28, 25] have revolutionized the field of visual generation, demonstrating remarkable capabilities that significantly outperform traditional approaches like Generative Adversarial Networks (GANs)[6] in terms of image quality and diversity.
-While these models excel at generating highly realistic imagery, a critical challenge persists: enabling precise and flexible control over the generation process to accommodate diverse and complex user requirements.
-Text-based conditioning has been a cornerstone in advancing controllable generation[28, 25, 13, 2, 23, 35], offering an intuitive interface for users to specify their desired outputs.
-However, text prompts alone often fail to convey precise spatial details and structural attributes that users wish to control.
-Consequently, recent research has explored complementary conditioning modalities for guiding diffusion models, with image-based control emerging as a particularly effective approach[41, 39, 22, 43, 15].
-This multi-modal conditioning strategy enables more detailed and accurate control over the generation process, addressing the limitations inherent in purely text-based interfaces.
-Current image conditioning methods can be broadly categorized into spatially aligned and non-spatially aligned approaches.
-Spatially aligned tasks such as sketch-to-image and inpainting require direct correspondence between conditioning and output images, typically achieved through methods like ControlNet[41] that inject conditioning features in a spatially-preserving manner.
-In contrast, non-spatially aligned applications including subject driven generation and style transfer, as demonstrated by IP-Adapter[39], often employ pre-trained encoders like CLIP[27] to extract global features for integration via cross-attention mechanisms.
-Despite the effectiveness of existing image-conditioned approaches, they present several limitations that hinder their efficiency and flexibility [41, 39, 22].
-First, most existing methods are designed specifically for UNet-based architectures [29, 41, 22, 39, 31, 42, 43, 40, 24, 16], as seen in Stable Diffusion models [28, 25].
-While these approaches work well with UNet’s encoder-decoder structure, they may not translate effectively to the more advanced Diffusion Transformer (DiT) models[23] that have demonstrated superior image generation quality[13, 2].
-Additionally, current approaches typically specialize in either spatially aligned [22, 43, 41] or non-spatially aligned tasks [39, 42, 17, 15, 12], lacking a unified architecture to handle both control types effectively.
-This specialization often requires practitioners to employ different methods for different control scenarios, increasing system complexity and implementation overhead.
-Furthermore, these methods rely heavily on additional network structures [39, 42, 22, 43, 41, 17], which introduce substantial parameter overhead.
-To address these limitations, we propose a parameter-efficient approach for incorporating image-based control into DiT architectures[23].
-Our method reuse the model’s existing VAE encoder[28] to process conditioning images.
-Following the same token processing pipeline as noisy image tokens, we augment the encoded features with learnable position embeddings[34], and seamlessly integrate them alongside latent noise in the denoising network.
-This design enables direct multi-modal attention interactions[23, 30] between condition and generation tokens throughout the DiT’s transformer blocks, facilitating efficient information exchange and control signal propagation.
+The ability to generate high-quality images with precise user control remains a central challenge in computer vision.
+While diffusion models [10, 28] have significantly advanced the state of the art in image generation, surpassing traditional GAN-based approaches [7] in both quality and diversity, the problem of fine-grained control persists.
+Text-conditioned diffusion models [28, 25, 5, 14] serve as the primary paradigm for controllable generation, enabling users to guide the synthesis process through natural language descriptions.
+However, these models fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need to control.
+To address this limitation, recent works have turned to image-based control methods [42, 41, 22, 44, 16, 34, 38, 17, 26, 37] as a promising solution to achieve fine-grained control.
+These approaches enable users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [28].
+However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules, leading to significant parameter overhead [42, 41].
+Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [42, 22, 44, 26]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [41, 17, 16, 34]), but rarely both.
+Third, current approaches are built primarily for UNet architectures [29].
+These approaches may yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [24] due to fundamental architectural differences, despite the latter’s superior generation capabilities [14, 2, 5].
+These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control.
+We propose OminiControl, an omni-capable yet minimal framework that achieves effective control by leveraging existing components within the DiT architecture.
+To achieve the goal of minimal architectural and parameters overhead, OminiControl first utilizes the built-in VAE encoder to project conditioning images into the same latent space as the target images.
+The conditional tokens are concatenated with the noisy latent and processed by the DiT using multi-modal attention [23, 30, 24].
+This unified sequence allows the denoising network to handle both signals directly, requiring only minimal LoRA [4] fine-tuning of the existing transformer blocks.
+Unlike existing methods such as IP-Adapter [41] that relies on separate CLIP [27] feature extractors and ControlNet [15] that requires additional control modules, OminiControl significantly reduces architectural complexity and parameter overhead.
+Building on this, OminiControl unifies controlled generation tasks using the DiT architecture.
+By modifying explicit position embeddings [35] in the attention mechanism, it handles both spatially aligned and non-aligned tasks.
+By explicitly controlling condition token positions in the attention mechanism, OminiControl not only handles non-aligned control tasks through semantic attention, but also achieves spatially-aligned control through flexible token interactions, rather than relying on rigid spatial feature addition as in ControlNet [42].
+Alongside the design of OminiControl, to fully realize its potential in diverse control scenarios, we address a critical challenge in subject-driven generation – the scarcity of high-quality training data for identity preservation.
+To this end, we develop an automated data synthesis pipeline specifically designed for subject-driven generation tasks.
+This pipeline builds upon a key observation that DiT models can generate identity-consistent image pairs with remarkable consistency by simply providing appropriate prompts [14].
+Using this pipeline, we create a comprehensive dataset Subjects200K including over 200,000 diverse images with rich identity variations.
+To facilitate future research, we have released both our dataset and the complete pipeline implementation as open-source resources [^2] Code and dataset are available at: https://github.com/Yuanshi9815/Subjects200K.
+Extensive experiments demonstrate that our approach achieves better performance across diverse control tasks.
+Building upon the state-of-the-art DiT model Flux [14], we evaluate OminiControl on a comprehensive suite of tasks including edge-guided generation, depth-guided generation, and subject-driven generation.
+Our method outperforms existing approaches such as ControlNet [42], T2I-Adapter [22], and IP-Adapter [41] in their original UNet implementations, as well as their community adaptations on the DiT architecture [15, 39].
+Notably, OminiControl achieves these superior results while maintaining remarkable efficiency – requiring only $0.1\%$ additional parameters compared to the base model.
+In summary, we highlight our contributions as follows:
+  • We propose OminiControl, a minimal and universal control framework for DiT models that: (1) minimizes architectural changes by reusing existing components , introducing only 0.1% additional parameters; and (2) achieves unified control by integrating condition tokens into the original multi-modal attention and leveraging position embeddings to handle both spatially-aligned and unaligned tasks.
+  • We develop an automated data synthesis pipeline that leverages DiT’s inherent capability in generating identity-consistent image pairs.
+Using this, we create and release Subjects200K, a large-scale dataset containing over 200,000 diverse images for subject-driven generation.
+  • We conduct extensive experiments on various control tasks including edge-guided, depth-guided, and subject-driven generation.
+Building upon Flux, OminiControl outperforms existing approaches in both their original UNet implementations and DiT adaptations.
 Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
-We implemented our method on the high-performing DiT-structured diffusion model, FLUX.1-dev[13], a large-scale model containing 12 billion parameters.
-Extensive experiments on edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation indicate that our DiT-based approach yields better results compared to both UNet-based implementations[41, 39, 7] and their community adaptations on the FLUX.1 model [14, 37].
-For identity-preserving generation, we developed a novel data synthesis pipeline that generates high-quality, identity-consistent image pairs.
-Using this pipeline, we created a comprehensive dataset comprising over 200,000 diverse images.
-To facilitate future research in this direction, we will release both our dataset and the complete pipeline implementation as open-source resources [^2] More details are available in the supplementary material..
-In summary, we highlight our contributions as follows:
-  1. 1.
-    We present a parameter-efficient method for enabling image-conditioned control in Diffusion Transformer (DiT) models, achieving both spatially aligned and non-spatially aligned control within a unified framework.
-  2. 2.
-    We demonstrate the effectiveness of our approach through extensive experiments across diverse control tasks, including edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation, consistently outperforming existing methods on both UNet implementations and their DiT adaptations.
-  3. 3.
-    We develop and release Subjects200K, a high-quality dataset of over 200,000 subject-consistent images, along with an efficient data synthesis pipeline, providing valuable resources to the research community for further exploration of subject-consistent generation tasks.
 2 Related works
-2.1 Diffusion-based models
-Diffusion-based methods have emerged as a powerful framework for image generation[9, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1].
+2.1 Diffusion models
+Diffusion models have emerged as a powerful framework for image generation[10, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1].
 Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28].
@@
 UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size.
-However, these methods rely on spatially adding condition information to the denoising network’s hidden states, inherently limiting their effectiveness for non-spatial tasks like subject-driven generation.
+However, these methods rely on spatially adding condition features to the denoising network’s hidden states, inherently limiting their effectiveness for spatially non-aligned tasks like subject-driven generation.
 IP-Adapter [41] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [43] further enhances identity preservation in image-conditioned tasks.
@@
 3.1 Preliminary
-The Diffusion Transformer (DiT) model [23], employed in architectures like FLUX.1 [13], Stable Diffusion 3 [28], and PixArt [2], uses a denoising network of transformer blocks to refine noisy image tokens iteratively.
-Each transformer block processes two types of tokens: noisy image tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2).
-These tokens are embedded into hidden states $X$ and $C_{\text{T}}$, which maintain consistent shapes throughout the transformer blocks.
-In each DiT block, after normalizing $X$ and $C_{\text{T}}$, they are processed by the core MM-Attention module [30], which employs Rotary Position Embedding (RoPE) [34] to incorporate positional dependencies across tokens.
-For a token at position $(i,j)$ in the 2D grid, RoPE applies rotation matrices to the query and key projections:
-$Q_{X}(i,j)$ $=W_{Q}\left(X_{i,j}\cdot R(i,j)\right),$ (1)
-$K_{X}(i,j)$ $=W_{K}\left(X_{i,j}\cdot R(i,j)\right),$ (2)
-where $R(i,j)$ is the rotation matrix at position $(i,j)$.
-Similarly, the text condition tokens $C_{\text{T}}$ have their query and key projections defined in the same way, with all text token positions set to $(0,0)$ in FLUX.1.
-After applying RoPE, the queries, keys, and values from both token types are concatenated to form unified matrices $Q_{Z}$, $K_{Z}$, and $V_{Z}$, representing the combined token set $Z=[X;C_{\text{T}}]$.
-The MM-Attention operation is then computed as:
-$\text{MMAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}\right)V_{Z},$ (3)
-enabling interactions between image and text condition tokens through the attention mechanism.
+The Diffusion Transformer (DiT) model [24], employed in architectures like FLUX.1 [14], Stable Diffusion 3 [28], and PixArt [2], uses transformer as denoising network to refine noisy image tokens iteratively.
+A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2).
+Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
+In FLUX.1, each DiT block consists of layer normalization followed by Multi-Modal Attention (MMA) [23], which incorporates Rotary Position Embedding (RoPE) [35] to encode spatial information.
+For image tokens ${X}$, RoPE applies rotation matrices based on the token’s position $(i,j)$ in the 2D grid:
+${X}_{i,j}\rightarrow{X}_{i,j}\cdot R(i,j),$ (1)
+where $R(i,j)$ is the rotation matrix at position $(i,j)$. Text tokens ${C}_{\text{T}}$ undergo the same transformation with their positions set to $(0,0)$.
+The multi-modal attention mechanism then projects the position-encoded tokens into query $Q$, key $K$, and value $V$ representations. It enables the computation of attention between all tokens:
+$\text{MMA}([{X};{C}_{\text{T}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
+where $[{X};{C}_{\text{T}}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
 Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
-3.2 Image condition integration
-Our approach first encodes the condition image through the model’s VAE, projecting it into the same latent space as the noisy image tokens to form $\mathbf{C}_{\text{I}}\in\mathbb{R}^{N\times d}$.
-Previous methods like ControlNet [41] and T2I-Adapter [22] incorporate the condition image by spatially aligning and adding its hidden states directly to those of the noisy image tokens:
-$H_{X}=H_{X}+H_{C_{\text{I}}},$ (4)
-where $H_{X}$ represents the combined hidden states for further processing, with $H_{C_{\text{I}}}$ being the hidden states from the condition image.
-While this approach proves effective for spatially aligned tasks, it faces two key limitations: (1) it lacks flexibility when handling non-aligned scenarios, and (2) even in spatially aligned cases, the direct addition of hidden states constrains token interactions, potentially limiting the model’s performance.
+3.2 OminiControl
 (a) Training losses for different image condition integration methods.
@@
 Figure 4: Training loss comparisons.
-In contrast, to enable non-aligned control tasks and provide greater conditioning flexibility, our method processes condition image tokens uniformly with text and noisy image tokens, integrating them into a unified sequence:
-$Z=[X;C_{\text{T}};C_{\text{I}}],$ (5)
-where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$.
-This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2).
+Building upon the DiT architecture, with FLUX.1 [14] as our implementation base,we aim to develop OminiControl, an omni-capable yet minimal control framework that guides generation through visual conditions.
+Leveraging existing architecture
+To achieve both omni-capability and minimal architectural changes, OminiControl first reuses the VAE [12, 28] encoder from the base DiT model to project conditioning images into the same latent space as noisy image tokens.
+This approach contrasts with previous methods that rely on separate feature extractors like CLIP [27, 41] or additional control modules [43], significantly reducing architectural complexity.
+The encoded condition tokens ${C}_{\text{I}}$ share the same dimensionality and latent space as the noisy image tokens ${X}$, enabling them to be directly processed by the transformer blocks.
+Since both condition and image tokens reside in the same latent space, OminiControl leverages the existing DiT blocks to process them jointly, requiring only minimal fine-tuning through LoRA [4] to adapt these blocks for handling the additional condition tokens, instead of duplicating the entire network as done in ControlNet [42].
+Unified sequence processing
+Building upon these shared latent space, we next need to integrate condition tokens into the model for flexible control.
+Previous methods like ControlNet [42] and T2I-Adapter [22] incorporate condition images through direct feature addition:
+${X}\leftarrow{X}+{C}_{\text{I}},$ (3)
+where the condition features ${C}_{\text{I}}$ are spatially aligned and added to the noisy image tokens ${X}$.
+While effective for spatially aligned tasks, this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
+Instead, OminiControl directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing.
+This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
+As shown in Figure 5, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns.
+Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature addition (Figure 4(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
 Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$.
@@
 Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
-The comparative result shows that our approach achieves higher generation quality and better alignment with the conditions compared to the direct adding method, as illustrated in Figure 3.
-Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the direct adding approach.
-Besides, the effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
-3.3 Adaptive position embedding
-Our unified sequence design allows for flexible integration of condition image tokens, but this requires incorporating positional information to ensure effective interaction with noisy image tokens.
-The relative positioning of these tokens is critical, as it directly affects the model’s learning efficiency and generalization capability.
-In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information.
-For a 512$\times$512 target image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token is assigned a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$.
-This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
-For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the noisy image.
-However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)).
-Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
-3.4 Condition strength factor
-The unified attention mechanism we adopt not only enables flexible token interaction but also allows us to precisely control the influence of condition images.
-Specifically, we designed a method that allows for manual adjustment of the condition image’s effect during inference.
-For a given strength factor $\gamma$, setting $\gamma=0$ removes the condition image’s influence, resulting in an output based purely on the original input.
-At $\gamma=1$, the output fully reflects the condition image, and as $\gamma$ increases beyond 1, the condition’s effect becomes even more pronounced.
-To achieve this controllability, we introduce a bias term into the original MM-Attention operation. Specifically, we modify Equation 3 to:
-$\text{BiasedAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}+\text{bias}(\gamma)\right)V_{Z},$ (6)
-[…]
-$\text{bias}(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (7)
-This design ensures that the strength factor $\gamma$ only affects the attention weights between noisy image tokens and condition image tokens, while maintaining the original attention patterns for text tokens and within-modality interactions.
+Position-aware token interaction
+While the unified sequence approach allows flexible token interactions, encoding position information for the newly appended conditional tokens is not straightforward.
+In FLUX.1, the RoPE [35] mechanism requires each token to have an assigned position index $(i,j)$, which might directly influence how tokens attend to each other during multi-modal attention.
+Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$.
+Text tokens ${C}_{\text{T}}$ maintain fixed positions at $(0,0)$ following FLUX.1’s convention.
+Therefore, it is essential to assign proper position indexing the conditioned tokens.
+In our method, we assign position indices differently for spatially aligned and non-aligned tasks.
+For condition image tokens ${C}_{\text{I}}$, we propose a dynamic positioning strategy based on the control task:
+$(i,j)_{{C}_{\text{I}}}=\begin{cases}(i,j)_{{X}}&\text{for aligned tasks}\\ (i,j)+\Delta&\text{for non-aligned tasks}\end{cases}$ (4)
+where $\Delta$ is a fixed offset vector.
+For spatially aligned tasks like edge-guided generation, condition tokens share positions with their corresponding noisy image tokens, facilitating precise spatial correspondence.
+For non-aligned tasks like subject-driven generation, we shift condition tokens’ positions by $\Delta$, ensuring no spatial overlap with ${X}$.
+As demonstrated in Figure 4(b), this shifting leads to faster convergence and better performance for non-aligned tasks.
+Controllable conditioning strength
+While OminiControl’s unified sequence processing and multi-modal attention enable effective token interactions during training, practical applications often require adjustable conditioning strength at test time.
+We achieve this by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
+$\text{MMA}([{C}_{\text{T}};{X};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
+where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{C}_{\text{T}};{X};{C}_{\text{I}}]$.
+Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
+$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
+This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$.
+At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it.
+This allows users to flexibly adjust the balance between text and image conditions without retraining the model.
 3.3 Subjects200K datasets
@@
 Each pair of images shows the same object in varying positions, angles, and lighting conditions.
-The dataset includes diverse objects such as clothing, furniture, vehicles, and animals, totaling over 200,000 images.
-This dataset, along with the generation pipeline, will be publicly released.
-Training models for subject-consistent generation typically requires paired images that maintain identity consistency while exhibiting variations in pose, lighting, and other attributes.
-Previous methods like IP-Adapter [39] use identical images for conditioning and target pairs, which proves effective for their approaches.
-However, in our framework, this setup leads to overfitting, causing the model to generate outputs nearly identical to the inputs.
-To overcome these limitations, we developed a dataset featuring images that preserve subject identity while incorporating natural variations.
-While existing datasets [31, 12, 15, 17] address similar needs, they often face constraints in either quality or scale.
-We therefore propose a novel synthesis pipeline leveraging FLUX’s inherent capability to generate pairs of visually related images from carefully crafted prompts.
-Our pipeline utilizes ChatGPT-4o to generate over 20,000 diverse image descriptions, which guide FLUX in producing more than 200,000 images (Figure 6).
-The generated images undergo quality assessment using ChatGPT-4o’s vision capabilities, ensuring both consistency and diversity in the final dataset.
-The curated dataset and complete generation pipeline are publicly available [^3] More details are provided in the supplementary material.
-Dataset and code are available at https://github.com/Yuanshi9815/Subjects200K.
+The dataset includes diverse objects such as clothing, furniture, and animals, totaling over 200,000 images.
+A critical challenge in developing universal control frameworks lies in obtaining high-quality training data for subject-driven generation.
+While OminiControl’s architecture enables flexible control, effective training requires data that maintains subject consistency while incorporating natural variations in pose, lighting, and context.
+Existing solutions often rely on identical image pairs [41] or limited-scale datasets [31, 13], which present several limitations.
+Using identical pairs can lead to overfitting in our framework, causing the model to simply reproduce the input.
+Meanwhile, existing datasets with natural variations often lack either the scale or diversity needed for robust training.
+Condition Model Method Controllability General Quality Text Consistency
+[…]
+Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.304
+T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.305
+FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.192
+Ours 0.38 20.63 0.40 0.61 75.91 0.308
+Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.308
+T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.309
+FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.212
+Ours 903 27.26 0.39 0.55 75.06 0.308
+Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.293
+Ours 132 11.49 0.87 0.39 67.63 0.300
+Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.278
+Ours 24 10.23 0.73 0.43 70.74 0.301
+Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.300
+FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.306
+Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks.
+We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score).
+For F1-Score, higher is better; for MSE, lower is better.
+Best results are shown in bold.
+To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [15] can generate identity-consistent image pairs when provided with appropriate prompts.
+Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^3] Detailed prompt design and synthesization pipeline are provided in the Appendix A.:
+  • Prompt Generation.
+We use GPT-4o to generate over 30,000 diverse subject descriptions.
+Each description represents the same subject in multiple scenes.
+The goal is to highlight distinct objects or subjects while allowing natural variations.
+  • Paired-image synthesis.
+We reorganized the collected descriptions into structured prompts.
+Each prompt describes the same subject in two different scenes.
+The template of such prompt is shown in Supplementary Material S3.These prompts are then fed into FLUX to generate image pairs.
+Each pair is designed to maintain subject consistency while varying in pose, lighting, and context.
+  • Quality Assesment. Finally, we use GPT-4o to evaluate the generated pairs. Misaligned pairs are removed to ensure identity consistency and high image quality.
+The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories including clothing, furniture, vehicles, and animals.
+Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control.
+To facilitate future research, we release both the dataset and our complete generation pipeline [^4] Dataset and code available at: https://github.com/Yuanshi9815/Subjects200K..
 4 Experiment
@@
 In subject-driven generation tasks, we switch to FLUX.1-schnell as we observed it tend to produce better visual quality.
-Condition Model Method Controllability General Quality Text Consistency
-[…]
-Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.75
-T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.74
-FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.53
-Ours 0.38 20.63 0.40 0.61 75.91 0.76
-Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.726
-T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.72
-FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.54
-Ours 903 27.26 0.39 0.55 75.06 0.728
-Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.78
-Ours 132 11.49 0.87 0.39 67.63 0.87
-Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.85
-Ours 24 10.23 0.73 0.43 70.74 0.90
-Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.84
-FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.80
-Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks.
-We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score).
-For F1-Score, higher is better; for MSE, lower is better.
-Best results are shown in bold.
+Figure 7: Qualitative results comparing different methods.
+Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization.
+Right: Subject-driven generation with beverage can, shoes and robot toy.
+Our method demonstrates superior controllability and visual quality across all tasks.
 Implement details.
@@
 We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01.
+For spatially aligned tasks, we use text-to-image-2M[8] dataset with the last 300,000 images.
+For subject-driven generation, we utilize our proposed Subjects200K dataset.
 The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each).
@@
 Baselines.
-For spatially-aligned tasks, we compare our method with both the original ControlNet [41] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [14], the FLUX.1 implementation of ControlNet.
+For spatially aligned tasks, we compare our method with both the original ControlNet [42] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [15], the FLUX.1 implementation of ControlNet.
 For subject-driven generation, we compare with IP-Adapter [41], evaluating its implementations FLUX.1 [39].
@@
 We conducted evaluations on two datasets.
-For spatially-aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42.
+For spatially aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42.
 For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
+Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
 4.2 Main result
-Figure 7: Qualitative results comparing different methods.
-Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization.
-Right: Subject-driven generation with beverage can, shoes and robot toy.
-Our method demonstrates superior controllability and visual quality across all tasks.
 Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks.
@@
 As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
-Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
-Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines.
-Our method demonstrates superior performance, particularly in identity preservation and modification accuracy.
-Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%.
-The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing.
-These quantitative results are further corroborated by user studies presented in Appendix B.1.
-Comparative paramter efficiency.
-As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods.
-For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%).
-Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
 Methods Base model Parameters Ratio
@@
 For our method, we also report results when using the original VAE encoder from FLUX.1.
+Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines.
+Our method demonstrates superior performance, particularly in identity preservation and modification accuracy.
+Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%.
+The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing.
+These quantitative results are further corroborated by user studies presented in Appendix B.1.
+Comparative paramter efficiency.
+As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods.
+For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%).
+Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
 4.3 Empirical studies
@@
 Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$.
-Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like depth-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
+Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and spatially non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
 Figure 10: Comparison of models trained with different data.

### v6

@@
 Abstract
-In this paper, we introduce OminiControl, a highly versatile and parameter-efficient framework that integrates image conditions into pre-trained Diffusion Transformer (DiT) models.
-At its core, OminiControl leverages a parameter reuse mechanism, enabling the DiT to encode image conditions using itself as a powerful backbone and process them with its flexible multi-modal attention processors.
-Unlike existing methods, which rely heavily on additional encoder modules with complex architectures, OminiControl (1) effectively and efficiently incorporates injected image conditions with only 0.10.1~{}0.10.1additional parameters, and (2) addresses a wide range of image conditioning tasks in a unified manner, including subject-driven generation and spatially-aligned conditions such as edges, depth, and more.
-Remarkably, these capabilities are achieved by training on images generated by the DiT itself, which is particularly beneficial for subject-driven generation.
-Extensive evaluations demonstrate that OminiControl outperforms existing UNet-based and DiT-adapted models in both subject-driven and spatially-aligned conditional generation.
-Additionally, we release our training dataset, Subjects200K, a diverse collection of over 200,000 identity-consistent images, along with an efficient data synthesis pipeline to advance research in subject-consistent generation.
-[^1] Code and dataset are available at https://github.com/Yuanshi9815/OminiControl
+We present OminiControl, a novel approach that rethinks how image conditions are integrated into Diffusion Transformer (DiT) architectures.
+Current image conditioning methods either introduce substantial parameter overhead or handle only specific control tasks effectively, limiting their practical versatility.
+OminiControl addresses these limitations through three key innovations: (1) a minimal architectural design that leverages the DiT’s own VAE encoder and transformer blocks, requiring just 0.1% additional parameters; (2) a unified sequence processing strategy that combines condition tokens with image tokens for flexible token interactions; and (3) a dynamic position encoding mechanism that adapts to both spatially-aligned and non-aligned control tasks.
+Our extensive experiments show that this streamlined approach not only matches but surpasses the performance of specialized methods across multiple conditioning tasks.
+To overcome data limitations in subject-driven generation, we also introduce Subjects200K, a large-scale dataset of identity-consistent image pairs synthesized using DiT models themselves.
+This work demonstrates that effective image control can be achieved without architectural complexity, opening new possibilities for efficient and versatile image generation systems.
 Figure 1: Results of our OminiControl  on both subject-driven generation (top) and spatially-aligned tasks (bottom). The small images in red boxes show the input conditions.
@@
 The ability to generate high-quality images with precise user control remains a central challenge in computer vision.
-While diffusion models [10, 28] have significantly advanced the state of the art in image generation, surpassing traditional GAN-based approaches [7] in both quality and diversity, the problem of fine-grained control persists.
-Text-conditioned diffusion models [28, 25, 5, 14] serve as the primary paradigm for controllable generation, enabling users to guide the synthesis process through natural language descriptions.
-However, these models fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need to control.
-To address this limitation, recent works have turned to image-based control methods [42, 41, 22, 44, 16, 34, 38, 17, 26, 37] as a promising solution to achieve fine-grained control.
-These approaches enable users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [28].
-However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules, leading to significant parameter overhead [42, 41].
-Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [42, 22, 44, 26]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [41, 17, 16, 34]), but rarely both.
+While diffusion models [17, 44] have significantly advanced image generation, surpassing traditional GAN-based approaches [11] in both quality and diversity, fine-grained control remains problematic.
+Text-conditioned models [44, 41, 9, 24] serve as the primary controllable generation paradigm, yet fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need.
+To address this limitation, some works have explored image-based control methods [61, 60, 37, 63, 27, 50, 57, 28, 42, 56, 29, 52, 15, 2, 30, 18, 14, 26, 46, 57, 54], enabling users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [44].
+However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules [61, 60, 38].
+Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [61, 37, 63, 42, 29, 52, 15, 2, 30]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [60, 28, 27, 50, 18][14, 33]), but rarely both.
 Third, current approaches are built primarily for UNet architectures [45].
-These approaches may yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [24] due to fundamental architectural differences, despite the latter’s superior generation capabilities [14, 2, 5].
+These approaches yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [40] due to fundamental architectural differences, despite the latter’s superior generation capabilities [24, 4, 9].
 These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control.
-We propose OminiControl, an omni-capable yet minimal framework that achieves effective control by leveraging existing components within the DiT architecture.
-To achieve the goal of minimal architectural and parameters overhead, OminiControl first utilizes the built-in VAE encoder to project conditioning images into the same latent space as the target images.
-The conditional tokens are concatenated with the noisy latent and processed by the DiT using multi-modal attention [23, 30, 24].
-This unified sequence allows the denoising network to handle both signals directly, requiring only minimal LoRA [4] fine-tuning of the existing transformer blocks.
-Unlike existing methods such as IP-Adapter [41] that relies on separate CLIP [27] feature extractors and ControlNet [15] that requires additional control modules, OminiControl significantly reduces architectural complexity and parameter overhead.
-Building on this, OminiControl unifies controlled generation tasks using the DiT architecture.
-By modifying explicit position embeddings [35] in the attention mechanism, it handles both spatially aligned and non-aligned tasks.
-By explicitly controlling condition token positions in the attention mechanism, OminiControl not only handles non-aligned control tasks through semantic attention, but also achieves spatially-aligned control through flexible token interactions, rather than relying on rigid spatial feature addition as in ControlNet [42].
-Alongside the design of OminiControl, to fully realize its potential in diverse control scenarios, we address a critical challenge in subject-driven generation – the scarcity of high-quality training data for identity preservation.
-To this end, we develop an automated data synthesis pipeline specifically designed for subject-driven generation tasks.
-This pipeline builds upon a key observation that DiT models can generate identity-consistent image pairs with remarkable consistency by simply providing appropriate prompts [14].
-Using this pipeline, we create a comprehensive dataset Subjects200K including over 200,000 diverse images with rich identity variations.
-To facilitate future research, we have released both our dataset and the complete pipeline implementation as open-source resources [^2] Code and dataset are available at: https://github.com/Yuanshi9815/Subjects200K.
-Extensive experiments demonstrate that our approach achieves better performance across diverse control tasks.
-Building upon the state-of-the-art DiT model Flux [14], we evaluate OminiControl on a comprehensive suite of tasks including edge-guided generation, depth-guided generation, and subject-driven generation.
-Our method outperforms existing approaches such as ControlNet [42], T2I-Adapter [22], and IP-Adapter [41] in their original UNet implementations, as well as their community adaptations on the DiT architecture [15, 39].
-Notably, OminiControl achieves these superior results while maintaining remarkable efficiency – requiring only $0.1\%$ additional parameters compared to the base model.
+We propose OminiControl , an omni-capable yet minimal framework that achieves effective and flexible control.
+For minimal architecture, OminiControl employs a parameter reuse strategy that leverages DiT’s existing components—particularly its VAE encoder and transformer blocks—which already possess the necessary capabilities to process visual control signals.
+By reusing these components with minimal fine-tuning, our approach dramatically reduces parameter overhead while maintaining control effectiveness, requiring only 0.1% additional parameters compared to the base model.
+To achieve omni-capability across diverse tasks, OminiControl introduces a unified sequence processing approach that fundamentally differs from previous methods’ rigid feature addition.
+By directly concatenating condition tokens with noisy image tokens in a unified sequence, we allow the multi-modal attention mechanism to discover appropriate relationships between tokens—whether spatial or semantic—without imposing artificial constraints.
+This flexibility is crucial for handling both spatially-aligned tasks like edge-guided generation and non-aligned tasks like subject-driven generation within a single framework.
+Building on this unified approach, OminiControl incorporates a dynamic positioning strategy that adaptively assigns position indices based on the task type, enabling true omni-capability without task-specific architectural modifications.
+Furthermore, for practical flexibility, we introduce an attention bias mechanism that allows users to dynamically adjust the influence of image conditions at inference time, providing crucial control over the generation process.
+The effectiveness of our architecture depends on high-quality training data, particularly for subject-driven generation.
+Recognizing that state-of-the-art DiT models can generate remarkably consistent image pairs of the same subject, we develop an automated data synthesis pipeline that creates Subjects200K—a dataset of over 200,000 diverse, identity-consistent images that provides the rich training signals needed to fully realize OminiControl ’s potential.
 In summary, we highlight our contributions as follows:
-  • We propose OminiControl, a minimal and universal control framework for DiT models that: (1) minimizes architectural changes by reusing existing components , introducing only 0.1% additional parameters; and (2) achieves unified control by integrating condition tokens into the original multi-modal attention and leveraging position embeddings to handle both spatially-aligned and unaligned tasks.
-  • We develop an automated data synthesis pipeline that leverages DiT’s inherent capability in generating identity-consistent image pairs.
-Using this, we create and release Subjects200K, a large-scale dataset containing over 200,000 diverse images for subject-driven generation.
-  • We conduct extensive experiments on various control tasks including edge-guided, depth-guided, and subject-driven generation.
-Building upon Flux, OminiControl outperforms existing approaches in both their original UNet implementations and DiT adaptations.
-Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
+  • We propose OminiControl , a minimal and universal control framework for DiT models that requires only 0.1% additional parameters while effectively handling both spatially-aligned and non-aligned tasks, demonstrating that extensive architectural modifications are unnecessary for effective image conditioning.
+  • We identify two key technical innovations that enable omni-capability: (1) unified sequence processing that outperforms traditional feature addition, and (2) adaptive position encoding that strategically assigns position indices based on task requirements.
+  • We design a flexible attention bias mechanism that allows precise adjustment of conditioning strength at inference time, enhancing practical control within the multi-modal framework without compromising performance.
+  • We develop and release Subjects200K, a large-scale dataset containing over 200,000 identity-consistent images to advance future research.
 2 Related works
@@
 Building on these architectural innovations, FLUX[24] incorporates transformer-based design with flow matching objectives [31], achieving state-of-the-art generation performance.
-2.2 Controllable generation with diffusion models
+2.2 Controllable generation
 Controllable generation has been extensively studied in the context of diffusion models.
@@
 IP-Adapter [60] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [62] further enhances identity preservation in image-conditioned tasks.
-Despite these advances [16, 19, 6], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
+Despite these advances [27, 32, 10, 38, 47, 18, 23, 14, 33], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
 3 Methods
@@
 The Diffusion Transformer (DiT) model [40], employed in architectures like FLUX.1 [24], Stable Diffusion 3 [44], and PixArt [4], uses transformer as denoising network to refine noisy image tokens iteratively.
-A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2).
+A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2(a)).
 Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
@@
 where $[{X};{C}_{\text{T}}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
-Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
 3.2 OminiControl
+(a) Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
+(b) Comparison of results
+Figure 2: Exploration of different methods for integrating image conditions.
 (a) Training losses for different image condition integration methods.
-(b) Training loss for shared vs. shifting position.
+(b) Training loss for shared vs. shifting position. (subject-driven)
 Figure 3: Training loss comparisons.
-Building upon the DiT architecture, with FLUX.1 [14] as our implementation base,we aim to develop OminiControl, an omni-capable yet minimal control framework that guides generation through visual conditions.
-Leveraging existing architecture
-To achieve both omni-capability and minimal architectural changes, OminiControl first reuses the VAE [12, 28] encoder from the base DiT model to project conditioning images into the same latent space as noisy image tokens.
+Building upon the DiT architecture, with FLUX.1 [24] as our base model, we aim to develop a minimal, omni-capable and controllable generation framework that accepts flexible control signals.
+This vision leads to our OminiControl , which we describe in this section.
+3.2.1 Minimal architecture
+To minimize the extra architectural and parameter overhead, OminiControl adopts a parameter reuse strategy.
+As its name implies, we reuses the VAE [22, 44] encoder from the base DiT model encode the condition image.
+These images are projected into the same latent space as the noisy input tokens, ensuring compatibility without introducing new modules.
+Following this, OminiControl processes both noisy image tokens and condition tokens jointly through the original DiT blocks.
+To adapt the shared DiT blocks for handling condition tokens, only lightweight LoRA fine-tuning [8] is applied, avoiding costly full-parameter updates.
 This approach contrasts with previous methods that rely on separate feature extractors like CLIP [43, 60] or additional control modules [62], significantly reducing architectural complexity.
-The encoded condition tokens ${C}_{\text{I}}$ share the same dimensionality and latent space as the noisy image tokens ${X}$, enabling them to be directly processed by the transformer blocks.
-Since both condition and image tokens reside in the same latent space, OminiControl leverages the existing DiT blocks to process them jointly, requiring only minimal fine-tuning through LoRA [4] to adapt these blocks for handling the additional condition tokens, instead of duplicating the entire network as done in ControlNet [42].
-Unified sequence processing
-Building upon these shared latent space, we next need to integrate condition tokens into the model for flexible control.
-Previous methods like ControlNet [42] and T2I-Adapter [22] incorporate condition images through direct feature addition:
-${X}\leftarrow{X}+{C}_{\text{I}},$ (3)
-where the condition features ${C}_{\text{I}}$ are spatially aligned and added to the noisy image tokens ${X}$.
-While effective for spatially aligned tasks, this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
-Instead, OminiControl directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing.
+Meanwhile, LoRA fine-tuning ensures parameter efficiency compared to duplicating the entire network as done in ControlNet [61].
+3.2.2 Omni-capable token interaction
+To achieve effective control across diverse tasks, OminiControl needs to enable flexible interactions between condition tokens and image tokens.
+We address this challenge through two key mechanisms: unified sequence processing and position-aware token interaction.
+Unified sequence processing.
+Building upon the shared latent space, we now think how to integrate condition tokens into the model for flexible control across.
+Previous methods [61, 37] incorporate condition images through direct feature adding:
+${h_{X}}\leftarrow{h_{X}}+h_{{C}_{\text{I}}},$ (3)
+where the condition features $h_{{C}_{\text{I}}}$ are spatially aligned and added to the noisy image features $h_{X}$.
+We first implement this direct adding approach as illustrated in Figure 2(a).
+While the bare effectiveness is shown in Figure 2(b), this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
+We then explored the unified sequence processing to integrate condition tokens into the model.
+Specifically, this approach directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing.
 This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
 As shown in Figure 4, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns.
-Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature addition (Figure 4(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
-Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$.
+Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature adding (Figure 3(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
+Figure 4: (a) Attention maps for the Canny-to-image task (with setting from Figure 2(b)), showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$.
 Strong diagonal patterns indicate effective spatial alignment.
 (b) Subject-driven generation task, with input condition and output image.
-Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
-Position-aware token interaction
+(Prompt: This item is placed on a table with Christmas decorations around it.) Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
+Position-aware token interaction.
 While the unified sequence approach allows flexible token interactions, encoding position information for the newly appended conditional tokens is not straightforward.
-In FLUX.1, the RoPE [35] mechanism requires each token to have an assigned position index $(i,j)$, which might directly influence how tokens attend to each other during multi-modal attention.
+FLUX.1 adopts the RoPE mechanism to encode spatial information for both image and text tokens.
 Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$.
-Text tokens ${C}_{\text{T}}$ maintain fixed positions at $(0,0)$ following FLUX.1’s convention.
-Therefore, it is essential to assign proper position indexing the conditioned tokens.
-In our method, we assign position indices differently for spatially aligned and non-aligned tasks.
-For condition image tokens ${C}_{\text{I}}$, we propose a dynamic positioning strategy based on the control task:
+We first assign the same position indices to condition tokens as noisy image tokens, it works well for spatially aligned tasks like edge-guided generation.
+However, for non-aligned tasks like subject-driven generation, this shared position indexing can lead to suboptimal performance due to spatial overlap between condition and noisy image tokens.
+But if we shift the position indices of condition tokens by a fixed offset $\Delta$ (e.g., $(0,32)$), ensuring no spatial overlap with noisy image tokens, the training convergence accelerates significantly and final performance improves.
+(see Figure 3(b) and Figure 5).
+Figure 5: The results from models with shared and shifted position indices. Both models are fully trained with 15k iterations.
+This observation suggests that for spatially aligned tasks, a shared position indexing facilitates direct spatial correspondence between condition and image tokens, while for non-aligned tasks, shared position indexing can constrain the model’s ability to establish semantic relationships.
+Hence, we propose a dynamic positioning strategy based on the control task:
 $(i,j)_{{C}_{\text{I}}}=\begin{cases}(i,j)_{{X}}&\text{for aligned tasks}\\ (i,j)+\Delta&\text{for non-aligned tasks}\end{cases}$ (4)
-where $\Delta$ is a fixed offset vector.
-For spatially aligned tasks like edge-guided generation, condition tokens share positions with their corresponding noisy image tokens, facilitating precise spatial correspondence.
-For non-aligned tasks like subject-driven generation, we shift condition tokens’ positions by $\Delta$, ensuring no spatial overlap with ${X}$.
-As demonstrated in Figure 4(b), this shifting leads to faster convergence and better performance for non-aligned tasks.
-Controllable conditioning strength
-While OminiControl’s unified sequence processing and multi-modal attention enable effective token interactions during training, practical applications often require adjustable conditioning strength at test time.
-We achieve this by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
-$\text{MMA}([{C}_{\text{T}};{X};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
-where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{C}_{\text{T}};{X};{C}_{\text{I}}]$.
+which makes OminiControl truly omni-capable by adapting to both spatially aligned and non-aligned control tasks.
+3.2.3 Control with flexibility
+Although our unified sequence processing and position-aware token interaction mechanisms effectively enable joint attention between conditions and images, they also present a new challenge.
+Unlike previous methods [60, 61] that could simply scale condition features (e.g., $h_{X}\leftarrow h_{X}+\alpha\cdot h_{C_{I}}$ where $\alpha$ controls strength), the joint attention approach of OminiControl does not inherently support adjustable conditioning strength, which is crucial for practical applications where users need to balance text and image influences.
+To address this limitation, we design a flexible control mechanism by introducing a bias term into the multi-modal attention computation.
+Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
+$\text{MMA}([{X};{C}_{\text{T}};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
+where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$.
 Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
-$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
+$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\mathbf{0}_{N\times N}\\ \log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{M\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
 This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$.
-At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it.
-This allows users to flexibly adjust the balance between text and image conditions without retraining the model.
+At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it, which makes OminiControl more flexible.
+3.2.4 Comparison with concurrent works
+Several recent works also explore controllable image generation with DiT models.
+Some approaches [24, 34, 12] employ channel concatenation to integrate condition tokens, which offers less flexibility than our unified sequence processing.
+Others [3, 6] focus exclusively on specific tasks such as style transfer or subject-driven generation, limiting their flexibility and generality.
+In contrast, our approach combines dynamic positioning with a flexible control mechanism to provide a more comprehensive and adaptable control framework compared to existing methods.
+Figure 6: Examples from our Subjects200K dataset. Each pair of images shows the same object in varying contexts.
 3.3 Subjects200K  datasets
-Figure 6: Examples from our Subjects200Kdataset.
-Each pair of images shows the same object in varying positions, angles, and lighting conditions.
-The dataset includes diverse objects such as clothing, furniture, and animals, totaling over 200,000 images.
 A critical challenge in developing universal control frameworks lies in obtaining high-quality training data for subject-driven generation.
@@
 Meanwhile, existing datasets with natural variations often lack either the scale or diversity needed for robust training.
-Condition Model Method Controllability General Quality Text Consistency
-[…]
-Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.304
-T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.305
-FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.192
-Ours 0.38 20.63 0.40 0.61 75.91 0.308
-Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.308
-T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.309
-FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.212
-Ours 903 27.26 0.39 0.55 75.06 0.308
-Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.293
-Ours 132 11.49 0.87 0.39 67.63 0.300
-Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.278
-Ours 24 10.23 0.73 0.43 70.74 0.301
-Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.300
-FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.306
-Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks.
-We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score).
-For F1-Score, higher is better; for MSE, lower is better.
-Best results are shown in bold.
-To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [15] can generate identity-consistent image pairs when provided with appropriate prompts.
-Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^3] Detailed prompt design and synthesization pipeline are provided in the Appendix A.:
-  • Prompt Generation.
-We use GPT-4o to generate over 30,000 diverse subject descriptions.
-Each description represents the same subject in multiple scenes.
-The goal is to highlight distinct objects or subjects while allowing natural variations.
+To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [25] can generate identity-consistent image pairs when provided with appropriate prompts [20, 19].
+Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^1] Detailed prompt design and synthesization pipeline are provided in the Section A of supplementary material.:
+  • Prompt Generation. We use GPT-4o to generate over 30,000 diverse subject descriptions. Each description represents the same subject in multiple scenes.
   • Paired-image synthesis.
-We reorganized the collected descriptions into structured prompts.
+We then reorganized the collected descriptions into structured prompts.
 Each prompt describes the same subject in two different scenes.
-The template of such prompt is shown in Supplementary Material S3.These prompts are then fed into FLUX to generate image pairs.
-Each pair is designed to maintain subject consistency while varying in pose, lighting, and context.
+The template of such prompt is shown in the Figure S3 of supplementary material.These prompts are then fed into FLUX to generate image pairs.
+Each pair is designed to maintain subject consistency while varying in context.
   • Quality Assesment. Finally, we use GPT-4o to evaluate the generated pairs. Misaligned pairs are removed to ensure identity consistency and high image quality.
-The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories including clothing, furniture, vehicles, and animals.
+The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories.
 Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control.
-To facilitate future research, we release both the dataset and our complete generation pipeline [^4] Dataset and code available at: https://github.com/Yuanshi9815/Subjects200K..
+To facilitate future research, we release both the dataset and our complete generation pipeline [^2] Dataset and code available at supplementary material..
 4 Experiment
@@
 In subject-driven generation tasks, we switch to FLUX.1-schnell as we observed it tend to produce better visual quality.
-Figure 7: Qualitative results comparing different methods.
-Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization.
-Right: Subject-driven generation with beverage can, shoes and robot toy.
-Our method demonstrates superior controllability and visual quality across all tasks.
+Figure 7: Qualitative comparison.
+Left: Spatially aligned tasks.
+Right: Subject-driven generation with beverage can and shoes.
+Our method demonstrates superior controllability and quality across all tasks.
+(More results are provided in supplementary materials.)
 Implement details.
 Our method utilizes LoRA [8]for fine-tuning the base model with a default rank of 4.
-To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 when processing non-condition tokens by default.
+To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 by default when processing non-condition tokens.
 Training.
@@
 For subject-driven generation, we compare with IP-Adapter [60], evaluating its implementations FLUX.1 [58].
-Evaluation metrics.
+Additionally, we also with the official FLUX.1 Tools [24] implementation.
+Evaluation of spatially aligned tasks.
 We evaluate our model on both spatially aligned tasks and subject-driven generation.
 For spatially aligned tasks, we assess two aspects: generation quality and controllability.
-Generation quality is measured using FID [9], SSIM, MAN-IQA [40], and MUSIQ [11] for visual fidelity, along with CLIP Score [27] for semantic consistency.
+Generation quality is measured using FID [16], SSIM, CLIP-IQA[53], MAN-IQA [59], MUSIQ [21], and PSNR[7] for visual fidelity, along with CLIP Text and CLIP Image scores [43] for consistency.
 For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.).
+We use 5,000 images from COCO 2017 validation set, and resize them to 512$\times$512, then generate task-specific conditions and associated captions as prompts with a fixed seed of 42.
+Evaluation of subject-driven generation.
 For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation.
-Detailed evaluation methodology is presented in Appendix B.1.
-Evaluation protocol.
-We conducted evaluations on two datasets.
-For spatially aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42.
-For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
+Details are presented in the Section B.2 of supplementary material.
+We test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[47] dataset with 5 different seeds, using one selected image per subject as the condition.
 Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
@@
 Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [61] and T2I-Adapter [37], as well as FLUX.1-based ControlNetPro [25].
-In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [40], and MUSIQ [11] scores.
+In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [55], CLIP-IQA[53], MAN-IQA [59], MUSIQ [21] and PSNR[7] scores.
 For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [16] improve from 30.38 to 11.49 for deblurring.
-The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality.
+The CLIP-Text and CLIP-Image metrics [43] indicate that our method maintains high consistency across all tasks, suggesting effective preservation of semantic alignment and image alignment while achieving better control and visual quality.
 As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
@@
 IP-Adapter 918M $\sim$7.6%
+FLUX.1 Tools 612M $\sim$5.1%
 Ours FLUX.1 / 12B 14.5M / 48.7M w/ Encoder $\sim$0.1/ $\sim$0.4w/ Encoder
@@
 For our method, we also report results when using the original VAE encoder from FLUX.1.
+Task Methods / Setting Base Model Controllability Image Quality Alignment
+[…]
+Canny ControlNet SD1.5 0.35 18.74 0.36 0.65 0.45 67.81 10.27 0.305 0.752
+T2I Adapter 0.22 20.06 0.35 0.57 0.39 67.89 9.53 0.305 0.748
+ControlNet Pro FLUX.1 0.21 98.69 0.25 0.48 0.37 56.91 9.22 0.192 0.537
+Flux Tools 0.20 22.13 0.32 0.66 0.60 75.47 9.69 0.308 0.701
+Ours 0.50 24.20 0.45 0.66 0.62 74.87 11.34 0.305 0.785
+Depth ControlNet SD1.5 923 23.03 0.34 0.64 0.47 70.73 10.63 0.308 0.726
+T2I Adapter 1560 24.72 0.28 0.61 0.39 69.99 9.50 0.309 0.721
+ControlNet Pro FLUX.1 2958 62.20 0.26 0.55 0.39 66.85 9.38 0.212 0.547
+Flux Tools 767 24.56 0.32 0.68 0.59 75.30 10.15 0.308 0.715
+Ours 537 31.04 0.39 0.68 0.60 74.04 10.53 0.305 0.749
+Mask ControlNet SD1.5 7588 13.14 0.68 0.58 0.42 67.22 18.96 0.300 0.848
+Flux Tools FLUX.1 6610 11.40 0.73 0.56 0.45 68.92 18.37 0.305 0.874
+Ours 6351 10.20 0.78 0.59 0.49 70.78 19.59 0.305 0.892
+Colorization ControlNet Pro FLUX.1 994 30.38 0.75 0.40 0.31 54.38 16.23 0.279 0.781
+Ours 73 10.37 0.92 0.56 0.48 69.40 21.56 0.305 0.884
+Deblur ControlNet Pro FLUX.1 338 16.27 0.64 0.55 0.43 70.95 20.53 0.294 0.853
+Ours 62 18.89 0.58 0.59 0.54 70.98 21.82 0.301 0.856
+Table 2: Quantitative comparison with baseline methods on five spatially aligned tasks.
+We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), Image Quality (SSIM, FID, CLIP-IQA, MAN-IQA, MUSIQ, PSNR), and Alignment (CLIP Text and CLIP Image).
+For F1-Score (used in Canny to Image task), higher is better; for MSE, lower is better.
+Best results are shown in bold.The second-best results are highlighted with underlines.
 Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines.
@@
 The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing.
-These quantitative results are further corroborated by user studies presented in Appendix B.1.
-Comparative paramter efficiency.
-As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods.
+These quantitative results are further corroborated by user studies presented in supplementary material B.2.
+Paramter efficiency As shown in Table 1, our approach achieves remarkable parameter efficiency compared to existing methods.
 For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%).
 Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
-4.3 Empirical studies
-Effect of training data.
-Figure 9: Comparison of models trained with different data.
-The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
-For subject-driven generation, our model takes a reference image of a subject (e.g., a plush toy or an object) and a text description as input, aiming to generate novel images of the same subject following the text guidance while preserving its key characteristics.
-To validate the effectiveness of our Subjects200K dataset described in Section 3.3, we compare two training strategies for this task.
-The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images.
-The second approach utilizes our Subjects200K dataset.
-As shown in Figure 9, the model trained with data augmentation only learns to replicate the input conditions with minimal changes.
-In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose.
-Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction.
-In contrast, our Subjects200K-trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
-Condition strength analysis.
-We evaluate our condition strength control through qualitative experiments.
-Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$.
-Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and spatially non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
-Figure 10: Comparison of models trained with different data.
-The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
+Condition strength factor We evaluate our condition strength control mechanism (Section 3.2.3) through qualitative experiments.
+Figure 9 shows generated results with varying strength factor $\gamma$.
+Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
+Figure 9: Demonstration of the condition strength control.
+4.3 Ablation Studies
+To better understand the key factors that influence our model’s control capabilities, we conducted comprehensive ablation studies examining parameter efficiency, architectural decisions, and component contributions.
 Impact of LoRA rank.
 We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task.
-As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects: image quality (measured by FID and SSIM), condition control capability (measured by F1 Score), while maintaining competitive text-image consistency (measured by CLIP-Score).
-Notably, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment where it achieves the highest CLIP-Score of 0.765, showing the efficiency of our approach even with limited parameters.
+As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects.
+However, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment, showing the efficiency of our approach even with limited parameters.
 Study Setting FID $\downarrow$ SSIM $\uparrow$ F1 Score $\uparrow$ CLIP Score $\uparrow$
@@
 FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens.
-As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability over the generation process.
-This suggests that allowing the condition signals to influence the entire transformer stack is crucial for achieving the desired levels of control over the output.
-Notably, this finding indicates that the preview approaches[15, 39, 41, 42, 22] of inserting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
+As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability.
+This suggests that allowing the condition signals to influence the entire transformer stack is crucial for generate better results.
+Notably, this finding indicates that the preview approaches[25, 58, 60, 61, 37] of injecting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
+Figure 10: Ablation study of critical modules for conditional generation. Given the censored Mona Lisa image and text prompt (bottom left), we test removing LoRA from different components.
+Critical module analysis.
+To identify essential components for effective control, we conducted fine-grained ablation studies as shown in Figure 10.
+Results demonstrate that normalization layers and attention projections—particularly query ($W_{Q}$) and key ($W_{K}$)—are critical for maintaining control quality.
+Removing LoRA from these components significantly degrades conditional rendering, while value projections ($W_{V}$) have less impact.
+These insights reveal that control signals might propagate through normalization pathways and attention routing mechanisms rather than feature transformation processes.
 5 Conclusion
-OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules.
-Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation.
-Results confirm OminiControl’s scalability and effectiveness in diffusion models.
+OminiControl  offers parameter-efficient image-conditional control for DiT across diverse tasks using a unified token approach, requiring only 0.1% additional parameters.
+The Subjects200K dataset — featuring over 200,000 high-quality, subject-consistent images—further supports advancements in subject-driven generation, with experimental results confirming OminiControl ’s effectiveness in both spatially-aligned and non-aligned tasks.
+However, the unified sequence approach increases the total number of tokens processed through the network, potentially limiting computational efficiency during inference.
+Addressing this token efficiency challenge while maintaining our method’s control capabilities represents an important direction for future research in parameter-efficient conditional generation.
 6 Acknowledgment
@@
 Appendix B Additional experimental results
-B.1 Evaluation for subject-driven generation
+B.1 Effect of training data
+Figure S4: Comparison of models trained with different data.
+The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200K generates novel views while preserving identity.
+For subject-driven generation, our model takes a reference image of a subject (e.g., a plush toy or an object) and a text description as input, aiming to generate novel images of the same subject following the text guidance while preserving its key characteristics.
+To validate the effectiveness of our Subjects200K  dataset described in Section 3.3, we compare two training strategies for this task.
+The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images.
+The second approach utilizes our Subjects200K  dataset.
+As shown in Figure S4, the model trained with data augmentation only learns to replicate the input conditions with minimal changes.
+In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose.
+Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction.
+In contrast, our Subjects200K -trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
+B.2 Evaluation for subject-driven generation
 Framework and criteria.
@@
 The results shown in Figure S5 corroborate our quantitative findings, with our method achieving superior performance across all evaluation criteria.
-B.2 Additional generation results
+B.3 Additional generation results
 We showcase more generation results from our method.
@@
 Figure S6: More results on Dreambooth dataset.
+Figure S7: More results on Dreambooth dataset.
 Figure S8: More results on other subject-driven generation tasks.